# Unifying KV Cache Compression for Large Language Models with LeanKV

### v2

@@ 6 Implementation
We implement LeanKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python. We first outline the architecture of LeanKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
6.1 LeanKV Architecture
The architecture of LeanKV is illustrated in Figure 6. At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers. Each GPU hosts one worker, responsible for executing a segment of the model. LeanKV utilizes tensor parallelism [54], where each worker handles a partition of each matrix multiplication and a subset of the attention heads. Each worker includes a dedicated memory manager to oversee the KV cache for its assigned attention heads, using the method described in Section 5. Due to dynamic sparsity patterns, memory requirements vary per head for each request, leading to an irregular memory layout. Additionally, each worker incorporates an execution engine for model computation. To support differentiated KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel. After computing key and value vectors, the KV compressor is invoked to compress them following the policy described in Section 4, storing the results in the KV cache. The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
Figure 6: LeanKV architecture.
@@ 6.2 Efficient Attention Kernel
We develop a custom GPU attention kernel that efficiently supports differentiated KV cache compression, fully harnessing the reduction in memory access volume and translating it into accelerated performance.
@@
Overall, we assign each CUDA thread block to process a single attention head per sequence. However, mixed-precision quantization introduces potential load imbalance, as high-precision tokens demand more bandwidth, which is the bottleneck of attention, than low-precision ones. Thanks to unified pages that store tokens of the same precision together, we mitigate this issue by having thread warps iterate over high-precision pages first, followed by low-precision pages, ensuring that threads within a warp remain load-balanced. Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values. These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively. To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated as follows.
@@
For key processing, each warp handles one page at a time, with threads in the warp divided into groups responsible for distinct keys. Within a group, each thread fetches its assigned keys in a vectorized manner, performs dequantization on the fly, and computes a partial dot product between the keys and corresponding query. Using a straightforward layout such as [$F$, $N_{\text{tokens}}$] for keys in each page, where $F$ denotes the feature dimension length, would cause threads within a group to fetch non-contiguous elements strided by $N_{\text{tokens}}$ when parallelizing across the feature dimension. Alternatively, using [$N_{\text{tokens}}$, $F$] would cause thread groups to fetch non-contiguous elements strided by $F$ when parallelizing across tokens. Both layouts suffer inefficient strided memory accesses, leading to low bandwidth utilization. Instead, we organize the layout of keys as [ $\frac{F}{K_{\text{vec}}\times K_{\text{group}}}$, $N_{\text{tokens}}$, $K_{\text{group}}$, $K_{\text{vec}}$], where $K_{\text{vec}}$ denotes the vectorization factor and $K_{\text{group}}$ denotes the number of threads per group. During each execution step, each thread fetches $K_{\text{vec}}$ consecutive elements, threads within a group collectively fetch $K_{\text{group}}$ adjacent chunks, and groups within the warp access contiguous chunks along the $N_{\text{tokens}}$ dimension. As a result, the combined memory accesses of all threads in a warp are contiguous, enabling memory coalescing and maximizing bandwidth utilization.
@@
For value processing, tokens in a page are evenly distributed across thread groups in a warp; within each group, individual threads perform sum reductions over specific ranges of the feature dimension and save their accumulation results in registers. Once all groups finish, a tree reduction aggregates these partial results and produces the output. The number of registers required per thread is proportional to the feature dimension range assigned to it, while during key processing each thread only requires a single register to store the partial dot product. As a consequence, vectorization along the feature dimension, as used in key processing, is not suitable for value processing, because it would significantly increase register pressure by forcing each thread to handle a larger portion of the feature dimension. To better align with the computation pattern of reduction across tokens during value processing, we apply vectorization to the token dimension instead, which reduces register pressure and ensures more effective parallelization. We organize the layout of values as [ $\frac{F}{V_{\text{group}}}$, $\frac{N_{\text{tokens}}}{V_{\text{vec}}}$, $V_{\text{group}}$, $V_{\text{vec}}$] accordingly.
@@ 7 Evaluation
We first evaluate the effectiveness of LeanKV’s differentiated KV cache compression policy and analyze the sensitivity of different models to KV cache compression. Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by LeanKV.
@@ 7.1 Experiment Setup
We evaluate LeanKV on five models spanning three major families: Llama3-8B and 70B [60], Qwen2.5-7B and 32B [28], and the recently released thinking model QwQ-32B [58], which generates extended chains of thought and exhibits strong performance on complex reasoning tasks. Our evaluation covers a diverse set of mainstream benchmarks across multiple task domains, including general knowledge (MMLU [26] and MMLU-Pro [55]), mathematics (GSM8K [13] and MATH [34]), code generation (HumanEval+ [11, 39] and MBPP+ [7, 39]), and long-context understanding (LongBench [8]). In addition, we assess LeanKV’s performance on QwQ-32B using two particularly challenging benchmarks: AIME24 [3], a high-school level mathematics competition dataset, and GPQA [48], which evaluates graduate-level science reasoning. Throughput and latency are measured on NVIDIA L40 GPUs, each equipped with 48 GB of memory [45].
@@ 7.2 Differentiated KV Compression Policy
Evaluating Differentiated KV Quantization. Figure 7 shows the performance of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama-3 8B & 70B and Qwen-2.5 7B. To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys. Results show that K8V4 matches the accuracy of the FP16 baseline across all models and benchmarks. In contrast, its mirror configuration, K4V8, exhibits noticeable performance degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks. This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama-3 8B. Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama-3 8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy. These results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
Figure 7: Performance of differentiated KV quantization.
Figure 8: Performance of dynamic vs. static sparsity.
@@
Evaluating Dynamic Sparsity. We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [35] and H2O [69], which allocates an equal memory budget to all attention heads. Figure 8 reports results on Qwen2.5-7B and Llama3-8B across GSM8K and HumanEval+. The x-axis indicates the percentage of tokens pruned, while the y-axis shows the resulting task accuracy. Dynamic sparsity significantly outperforms static sparsity in Llama3-8B, maintaining full accuracy with 50% tokens pruned on GSM8k, and 10% tokens pruned for the more sensitive HumanEval+. For Qwen2.5-7B, dynamic sparsity also consistently surpasses static sparsity, achieving accuracy higher than 80% on GSM8k with 50% tokens pruned. While the improvement is less pronounced on HumanEval+, dynamic sparsity still yields higher accuracy at equivalent pruning ratios. In summary, per-head dynamic sparsity is superior to static sparsity by leveraging head-specific significance.
@@
Figure 9: Calibrating the high- and low-precision threshold parameters $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
@@
Parameter Calibration. As described in Section 4, LeanKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length. A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $[\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise. To calibrate the appropriate parameters $\alpha_{h}$ and $\alpha_{l}$, we utilize the training split of MATH [34]. We select MATH as the calibration dataset as it is comprised of complex mathematical reasoning tasks with high text information density, where even the removal of a single symbol can invalidate an argument. This makes MATH a highly suitable choice for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks. Additionally, MATH provides a dedicated training split, ensuring that parameter fitting does not involve testing data and mitigating potential overfitting.
@@
We profile a consistent range of parameters across all evaluated models, as depicted in Figure 9. The x-axis represents the profiled parameter values, while the y-axis shows the resulting accuracy on the calibration dataset. Specifically, we profile $\alpha_{h}$ within the integer range $[1,5]$, progressively relaxing the high-precision threshold beyond the theoretical average. This adjustment accounts for aggregating scores from multiple attention heads by maximum for GQA, as mentioned in Section 4. The profiled values of $\alpha_{h}$ achieve FP16-equivalent accuracy across all models, except for Qwen2.5-7B, which exhibits heightened sensitivity to 4-bit key quantization. Based on our profiling, we select the values of $\alpha_{h}$ that yield the highest accuracy for each model: $\alpha_{h}=1$ for Llama-3 8B and 70B, $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and $\alpha_{h}=0$ for Qwen2.5-7B. Subsequently, we profile the low-precision threshold $\alpha_{l}$, while keeping $\alpha_{h}$ fixed. We observe that setting $\alpha_{l}=0.1$ induces at least 5% accuracy degradation on the larger models. As a result, we profile five equally spaced values within the range $[0,0.1]$ and select the one yielding the highest accuracy for each model. The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 00 for the remaining models. While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, LeanKV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
@@
Figure 10: KV cache memory normalized to vLLM vs. benchmark performance tradeoff of LeanKV.
@@
Evaluating Differentiated Compression Policy. Finally, we evaluate the combined benefits of LeanKV’s KV cache compression policy. We compare LeanKV against several state-of-the-art baselines, including pruning-based methods such as H2O [69], SnapKV [35], and DuoAttention [63], quantization-based approaches such as 4-bit KV and KIVI [41], as well as KV partial loading techniques like Quest [56]. Table 1 summarizes the accuracy and KV cache memory usage of LeanKV normalized to vLLM [32] baseline, with the parameters tuned as described, for the evaluated non-thinking models. For the pruning-based baselines and Quest, we tune their memory usage to 50% of the baseline. LeanKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory. In contrast, all other baselines incur more significant accuracy degradation at comparable or higher memory usage. Furthermore, LeanKV dynamically adapts its memory usage based on the task’s information density. For example, LeanKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
@@
LeanKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench. For Llama3-8B and Qwen2.5-7B, LeanKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage. Notably, LeanKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
@@
Furthermore, the performance gap between LeanKV and baseline methods becomes even more pronounced for the thinking model QwQ-32B on more challenging benchmarks, as demonstrated in Table 3. LeanKV matches the FP16 baseline performance using only $27.4\%$ of the memory on average, while all other methods experience significant performance degradation. On the math competition benchmark AIME24, the most accurate baseline, 4-bit KV quantization incurs a $16.0\%$ accuracy degradation. On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero.
@@
Lastly, Figure 10 depicts the performance-memory trade-offs of LeanKV across representative benchmarks under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B. We also include the accuracy of baseline methods within similar memory usage ranges for comparison. LeanKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness. In contrast, baseline methods experience significant accuracy degradation, particularly pruning-based and KV partial loading methods, which suffer rapid accuracy declines as memory usage decreases. In summary, LeanKV provides a superior performance-memory tradeoff compared to state-of-the-art methods.
@@
Table 1: Accuracy and memory usage of LeanKV and baseline methods across models and benchmarks. Numbers in parentheses represent memory usage normalized to FP16. The memory usage is 50% for SnapKV, Quest, H2O and DuoAttn, 25% for INT4, and 12.5% for KIVI due to 2-bit quantization.
@@
Table 2: Evaluating LeanKV on LongBench. LeanKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
@@
Table 3: Evaluating LeanKV on QwQ-32B thinking model.

### v3

@@ 6 Implementation
We implement DiffKV on top of vLLM, comprising 4.5K lines of CUDA/C++ code and 9K lines of Python. We first outline the architecture of DiffKV and then detail our custom GPU attention kernel, designed to efficiently support the differentiated KV cache compression.
6.1 DiffKV Architecture
The architecture of DiffKV is illustrated in Figure 7. At each inference step, the scheduler batches as many requests as possible within the available GPU memory to maximize throughput, sending the selected requests to all workers. Each GPU hosts one worker, responsible for executing a partition of the model [63, 33, 52, 53]. To facilitate efficient execution, each worker includes a dedicated memory manager that oversees the KV cache for its assigned attention heads, as outlined in Section 5, and an execution engine for model computation. To support differentiated KV cache compression, the execution engine integrates a KV compressor and a custom GPU attention kernel. After computing key and value vectors, the KV compressor is invoked to compress them following the policy described in Section 4, storing the results in the KV cache. The custom GPU attention kernel then efficiently computes the attention output using the compressed KV cache.
Figure 7: DiffKV architecture.
@@ 6.2 Efficient Attention Kernel
We develop a custom attention kernel that efficiently supports differentiated KV cache compression, leveraging reduced memory access volume to accelerate performance. Notably, the kernel design is portable across GPUs and other NPUs, as mainstream architectures support vectorized memory access and layout-aware tiling [37, 79]. We detail the GPU implementation in the following paragraphs.
@@
Overall, we assign each thread block to process a single attention head per sequence. To mitigate the potential load imbalance caused by mixed-precision quantization, we let thread warps iterate over high-precision pages first, followed by low-precision pages. Each page is processed in two phases: dot product between the query and keys to derive attention scores, and weighted sum of values. These two phases follow different computation patterns: reduction over the feature dimension and token dimension, respectively. To ensure coalesced and vectorized memory accesses for both phases, we design tailored data layouts and parallelization strategies, elaborated as follows.
@@
For key processing, each warp handles one page at a time, with threads divided into groups responsible for distinct key vectors. Within a group, each thread fetches its assigned key elements in a vectorized manner, performs dequantization on the fly, and computes a partial dot product between the key elements and corresponding query elements. Using a straightforward layout such as [$F$, $N_{\text{tokens}}$] for keys, where $F$ denotes the feature dimension length, would cause threads within a group to fetch non-contiguous elements strided by $N_{\text{tokens}}$ when parallelizing across the feature dimension. Alternatively, using $N_{\text{tokens}}$, $F$] would cause thread groups to fetch non-contiguous elements strided by $F$ when parallelizing across tokens. Both layouts suffer inefficient strided memory accesses, leading to low bandwidth utilization. Instead, we organize the layout of keys as [ $\frac{F}{K_{\text{vec}}\times K_{\text{group}}}$, $N_{\text{tokens}}$, $K_{\text{group}}$, $K_{\text{vec}}$], where $K_{\text{vec}}$ denotes the vectorization factor and $K_{\text{group}}$ denotes the number of threads per group. During each execution step, each thread fetches $K_{\text{vec}}$ consecutive elements, threads within a group collectively fetch $K_{\text{group}}$ adjacent chunks, and groups within the warp access contiguous chunks along the $N_{\text{tokens}}$ dimension. As a result, the combined memory accesses of all threads in a warp are contiguous, enabling memory coalescing and maximizing bandwidth utilization.
@@
For value processing, tokens in a page are evenly distributed across thread groups in a warp; within each group, individual threads perform sum reductions over the feature dimension and save their accumulation results in registers. Once all groups finish, a tree reduction aggregates these partial results and produces the output. The number of registers required per thread is proportional to the feature dimension range assigned to it, while during key processing each thread only requires a single register to store the partial dot product. Consequently, vectorization along the feature dimension, as used in key processing, is not suitable for value processing, because it would significantly increase register pressure by forcing each thread to handle a larger portion of the feature dimension. To better align with the computation pattern of reduction across tokens, we apply vectorization to the token dimension instead, which reduces register pressure and ensures more effective parallelization. We organize the layout of values as [ $\frac{F}{V_{\text{group}}}$, $\frac{N_{\text{tokens}}}{V_{\text{vec}}}$, $V_{\text{group}}$, $V_{\text{vec}}$] accordingly.
Additionally, for ultra-long sequences, the attention kernel supports parallelization along the sequence dimension. It splits the sequence into multiple segments, processes them in parallel using the aforementioned method within separate thread blocks, and then merges the results with minimal computation overhead.
@@ 7 Evaluation
We first evaluate the effectiveness of DiffKV’s differentiated KV cache compression policy. Next, we evaluate the efficiency of the memory manager as well as the end-to-end throughput improvements achieved by DiffKV.
@@ 7.1 Experiment Setup
We evaluate DiffKV on several models spanning three major families: Llama3-8B and 70B [68], Qwen2.5-7B and 32B [34], and the recent thinking models QwQ-32B [67], R1-Distill-Qwen-14B and R1-Distill-Llama-8B [26], which generate extended chains of thought and exhibit strong capabilities on complex reasoning tasks.
We select benchmarks that align with those commonly used in recent LLM technical reports [20, 34, 27, 45, 67], covering a diverse set of task domains, including general knowledge (MMLU [31] and MMLU-Pro [64]), mathematics (GSM8K [15] and MATH [41]), code generation (HumanEval+ [12, 46] and MBPP+ [8, 46]), and long-context understanding (LongBench [9]). In addition, we assess DiffKV’s generation quality on thinking models using two particularly challenging benchmarks: AIME24 [3], a mathematics competition dataset, and GPQA [57], which evaluates graduate-level science reasoning. We measure the throughput and latency on NVIDIA L40 GPUs, each with 48 GB of memory [54]. In all experiments, model weights are stored in FP16 precision.
@@ 7.2 Differentiated KV Compression Policy
Evaluating Differentiated KV Quantization. Figure 8 shows the model generation quality of differentiated KV quantization on GSM8K and HumanEval+, namely K8V4 and K4V2, compared to FP16 baselines, across Llama3-8B & 70B and Qwen2.5-7B. To validate our intuition that keys have a more significant impact than values (Section 3.1), we additionally evaluate the mirror configurations, namely K4V8 and K2V4, where values are stored with higher precision than keys, as well as more skewed variants, K8V2 and K4V1.
@@
Results show that K8V4 matches the accuracy of the FP16 baseline across all models and benchmarks. In contrast, its mirror configuration, K4V8, exhibits noticeable accuracy degradation, particularly for Qwen2.5-7B, where accuracy drops to nearly zero on both tasks. This pronounced sensitivity to 4-bit key quantization in Qwen2.5-7B is likely due to its GQA architecture, which applies aggressive KV compression with a queries-per-KV ratio of 7, substantially higher than the ratio of 4 used in Llama3-8B. Similarly, K4V2 retains over 65% of FP16 accuracy across both benchmarks on Llama3-8B and 70B, while the mirror configuration K2V4 results in near-zero accuracy. The more skewed variant K8V2 retains at least 83% of FP16 accuracy across all models, even for Qwen2.5-7B, reinforcing the importance of assigning higher precision to keys. On the other hand, K4V1 yields almost zero accuracy, suggesting that 2 bits is the lower bound for effective value quantization. In summary, these results confirm that keys play a more critical role than values, and demonstrate the effectiveness of differentiated KV quantization in preserving accuracy while enabling aggressive compression.
The accuracy trends across various differentiated KV quantization configurations motivate the design of adopting two precision levels: K8V4 and K4V2. K8V4 matches the generation quality of the FP16 baseline and is applied to important tokens to preserve model accuracy. K8V2 and K4V2, when applied uniformly, both lead to noticeable accuracy degradation. Among them, K4V2 offers better efficiency and is thus selected for compressing less significant tokens, enabling opportunistic KV cache savings with minimal quality loss. As we demonstrate later, the K8V4–K4V2 scheme achieves near-lossless generation quality across multiple models and benchmarks, obviating the need for an intermediate precision level like K8V2. Likewise, introducing an even lower precision level such as K4V1 is suboptimal: not only does K4V1 significantly degrade accuracy, but the additional memory savings from a three-level configuration K8V4–K4V2–K4V1 are marginal compared to the K8V4–K4V2 design. That said, future models employing more aggressive training-time KV compression (e.g., architectures with higher queries-per-KV ratios in GQA) may benefit from an additional high-precision level (e.g., FP16–K8V4 or FP16–K8V4–K4V2), which we leave to future investigation.
Figure 8: Accuracy of differentiated KV quantization.
Figure 9: Accuracy of dynamic vs. static sparsity.
@@
Evaluating Dynamic Sparsity. We evaluate the effectiveness of per-head dynamic sparsity (Section 3.3) in identifying critical tokens, comparing it to the static sparsity method used in SnapKV [42], which allocates an equal memory budget to all attention heads. Figure 9 reports results on Qwen2.5-7B and Llama3-8B across GSM8K and HumanEval+. The x-axis indicates the percentage of tokens pruned, while the y-axis shows the resulting task accuracy. Dynamic sparsity significantly outperforms static sparsity in Llama3-8B, maintaining full accuracy with 50% tokens pruned on GSM8K, and 10% tokens pruned for the more sensitive HumanEval+. For Qwen2.5-7B, dynamic sparsity also consistently surpasses static sparsity, achieving accuracy higher than 80% on GSM8K with 50% tokens pruned. While the improvement is less pronounced on HumanEval+, dynamic sparsity still yields higher accuracy at equivalent pruning ratios. In summary, per-head dynamic sparsity is superior to static sparsity by leveraging head-specific significance.
@@
Figure 10: Calibrating the high- and low-precision thresholds $\alpha_{h}$ and $\alpha_{l}$ on the training split of the MATH dataset.
@@
Parameter Calibration. As described in Section 4, DiffKV compares the significance score of a token with its theoretical average, $\frac{1}{N}$, where $N$ denotes the sequence length. A token is quantized to high precision if its score exceeds $\frac{\alpha_{h}}{N}$, to low precision if its score falls within the interval $\frac{\alpha_{l}}{N},\frac{\alpha_{h}}{N}]$, and is pruned otherwise. To calibrate the appropriate parameters $\alpha_{h}$ and $\alpha_{l}$, we utilize the training split of MATH [41]. We select MATH as the calibration dataset as it is comprised of complex mathematical reasoning tasks with high text information density, where even the removal of a single symbol can invalidate an argument. This makes MATH highly suitable for parameter fitting, as the resulting compression parameters are expected to generalize well across other tasks. Additionally, MATH provides a dedicated training split, ensuring that parameter fitting does not involve testing data and mitigating potential overfitting.
@@
We profile a consistent range of parameters across all evaluated models, as depicted in Figure 10. The x-axis represents the profiled parameter values, while the y-axis shows the resulting accuracy on the calibration dataset. Specifically, we profile $\alpha_{h}$ within the integer range $[1,5]$, progressively relaxing the high-precision threshold beyond the theoretical average. This adjustment accounts for aggregating scores from multiple attention heads by maximum for GQA, as mentioned in Section 4. The profiled values of $\alpha_{h}$ achieve FP16-equivalent accuracy across all models, except for Qwen2.5-7B, which exhibits heightened sensitivity to 4-bit key quantization. Accordingly, we select the values of $\alpha_{h}$ that yield the highest accuracy based on profiling results: $\alpha_{h}=1$ for Llama3-8B and 70B, and $\alpha_{h}=3$ for Qwen2.5-32B and QwQ-32B, and disable low-precision quantization for Qwen2.5-7B. Subsequently, we profile the low-precision threshold $\alpha_{l}$, while keeping $\alpha_{h}$ fixed. We observe that setting $\alpha_{l}=0.1$ induces at least 5% accuracy degradation on the larger models. As a result, we profile five equally spaced values within the range $[0,0.1]$ and select the one yielding the highest accuracy for each model. The final selected values of $\alpha_{l}$ are $0.02$ for Llama3-8B, $0.04$ for Qwen2.5-7B, and 0 for the remaining models. While the profiled values cover a limited portion of the entire parameter space, as will be demonstrated in the following sections, Diff-KV significantly outperforms existing baselines with these fitted parameters, highlighting its potential effectiveness.
@@
Figure 11: KV cache memory normalized to vLLM vs. benchmark accuracy tradeoff of DiffKV.
@@
Evaluating Differentiated Compression Policy. Finally, we evaluate the combined benefits of DiffKV’s KV cache compression policy. We compare DiffKV against several state-of-the-art baselines, including pruning-based methods such as H2O [76], SnapKV [42], and DuoAttention [70], quantization-based approaches such as 4-bit KV, KIVI [48], and QAQ [19], as well as KV partial loading techniques like Quest [65]. Table 1 compares the accuracy and KV cache memory usage of DiffKV and the best-performing baselines, both normalized to vLLM [39], for the evaluated non-thinking models. The parameters of DiffKV are tuned as described previously. For the pruning-based baselines and Quest, we tune their memory usage to 50% of the baseline. DiffKV achieves near-lossless accuracy relative to the FP16 baseline, with an average degradation of only $0.3\%$, while using $19.3\%$ to $36.7\%$ memory. In contrast, all other baselines incur more significant accuracy degradation at comparable or higher memory usage. Furthermore, DiffKV dynamically adapts its memory usage based on the task’s information density. For example, DiffKV allocates less memory for the 5-shot MMLU benchmark compared to the 0-shot HumanEval+, where the prompt consists solely of the function name and comments.
@@
DiffKV also outperforms baseline methods in long-context scenarios, as shown in Table 2, where we present results from one benchmark in each category of LongBench. For Llama3-8B and Qwen2.5-7B, DiffKV achieves memory usage of $14.9\%$ and $27.0\%$, respectively, while baseline methods are tuned to $25\%$ memory usage. Notably, DiffKV consistently achieves superior accuracy across all categories, with an average degradation of only $0.4\%$, while utilizing less or comparable memory compared to the baseline methods.
@@
Furthermore, the generation quality gap between DiffKV and baseline methods becomes even more pronounced for the thinking models QwQ-32B, R1-Distill-Qwen-14B and R1-Distill-Llama-8B on the more challenging benchmarks, as demonstrated in Table 3. DiffKV achieves FP16-comparable generation quality using $27.4\%$, $29.4\%$ and $23.5\%$ of the memory on average for QwQ-32B, R1-Distill-Qwen-14B and R1-Distill-Llama-8B, respectively, while all other methods experience significant generation quality degradation. On the math competition benchmark AIME24, the most accurate baseline, 4-bit KV quantization incurs a $16.0\%$ accuracy degradation. On the graduate-level science benchmark GPQA, the accuracy of pruning-based and KV partial loading methods drops to almost zero, while using $50\%$ of the full KV cache. The larger gap between DiffKV and prior work in long CoT generation highlights the increased challenges such tasks pose for KV cache compression. In long CoT generation, tokens are generated autoregressively based on previous tokens, and errors introduced by compression are accumulated and propagated across generation steps. In contrast, in long prompt scenarios, the majority of the text is provided as ground truth in the prompt, and the model-generated portion is significantly shorter, limiting the impact of errors introduced by KV cache compression.
@@
Lastly, Figure 11 depicts the accuracy-memory trade-offs of DiffKV under the profiled parameter values, namely $\alpha_{h}\in[0,5]$ (step size of 1) for all models and $\alpha_{l}\in[0.02,0.1]$ (step size of 0.02) for Llama3-8B and Qwen2.5-7B. We also include the accuracy of baseline methods within similar memory usage ranges for comparison. DiffKV consistently matches the FP16 baseline accuracy across different models and benchmarks within the profiled range of memory usage, demonstrating its robustness. In contrast, baseline methods experience significant accuracy degradation, particularly pruning-based and KV partial loading methods, which suffer rapid accuracy declines as memory usage decreases. In summary, DiffKV provides a superior accuracy-memory tradeoff compared to state-of-the-art methods.
@@
Table 1: Accuracy and memory usage of DiffKV and the best-performing baseline methods across models and benchmarks. Numbers in parentheses represent memory usage normalized to FP16. The memory usage is 50.0% for SnapKV, Quest and DuoAttn, 25.0% for INT4, 18.8% for QAQ, and 12.5% for KIVI. DiffKV achieves accuracy close to FP16, consistently outperforming prior quantization and pruning approaches.
@@
Table 2: Evaluating DiffKV on LongBench. DiffKV uses 14.9% memory for Llama3.1-8B and 27.0% for Qwen2.5-7B. For Quest and SnapKV, the memory usage is 25%.
@@
Table 3: Evaluating DiffKV on thinking models.
@@
Qwen-14B
[…]
Llama-8B
[…]
KV Compression Breakdown. Figure 12 shows the fraction of tokens that are pruned, quantized to low precision, and quantized to high precision across three benchmarks (MMLU, HumanEval+, and MATH) for three different models. Two key observations emerge. First, sparsity levels vary substantially across workloads. The general knowledge benchmark MMLU exhibits the highest sparsity, likely due to the lower information density of natural language and the use of 5-shot prompts. In contrast, the reasoning-intensive benchmarks MATH and HumanEval+ display lower sparsity, with HumanEval+ being the least sparse, partly due to its 0-shot setting. Second, despite being calibrated on challenging reasoning-heavy workloads, DiffKV’s KV compression policy adapts effectively to the varying sparsity patterns across different tasks, highlighting its generalizability to diverse scenarios.
Figure 12: Fraction of tokens pruned, quantized to low precision, and quantized to high precision across different benchmarks and models.